OpenStudy (anonymous):

1. For the f(x) = 5x + 8 and g(x) = x – 3. Calculate (f o g)(x)

OpenStudy (anonymous):

(f o g)(x) means f(g(x)) so, (f o g)(x) = f(g(x)) = f( x – 3) = ? u wanna try this step? plug in (x-3) everywhere you see the 'x' term in the f(x) function of 5x+8

OpenStudy (anonymous):

I came up with 23

OpenStudy (anonymous):

(f o g)(x) = = f(g(x)) = f( x – 3) = 5(x-3)+8 = 5x-5(3) + 8 = 5x - 7

OpenStudy (anonymous):

thank you!!!!!!!

OpenStudy (anonymous):

can you help me on this 1 to? 2. Consider the function. F(x) = (1/5)x - 4 a.) Find the inverse of f(x) and name it g(x). Show and explain your work.

OpenStudy (anonymous):

ur welcome ^_^

OpenStudy (anonymous):

to find an inverse means to rename F(x) to y then swap places of x and y, then solve for y then rename y to g(x)

OpenStudy (anonymous):

ok im about to try it out

OpenStudy (anonymous):

u can do it!! ^_^ haha

OpenStudy (anonymous):

im stuck on this problem. I came up with y= (5/1) but I don't think that's right

OpenStudy (anonymous):

F(x) = (1/5)x - 4 y = (1/5)x - 4 x = (1/5)y - 4 x+4 = (1/5)y 5(x+4)=y y = 5x + 20